OpenStudy (anonymous):

Multiply the following polynomials, then place the answer in the proper location on the grid. Write the answer in descending powers of a.(a - 5)(a 2 + 5a + 25)

OpenStudy (anonymous):

first multiply everything by \(a\) which means kick up the exponent by 1 you get \[a(x^2+5a+25)=a^3+5a^2+25a\] then multiply everything by \(-5\) you get \[-5(a^2+5a+25)=-5a^2-25a-125\] then combing like terms

OpenStudy (anonymous):

if you do this with pencil and paper you can write one on top of the other so the like terms are in a column like this a^3 + 5a^2 +25 a -5a^2 - 25 a -125 _______________________ a^3 -125

OpenStudy (anonymous):

answer is \(a^3-125\) and this is how you factor the difference of two cubes

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yw